# Lipschitz-Based Robustness Certification Under Floating-Point Execution

### 06 Mar 2026 v1

@@ 1. Introduction
Robustness is an important property for helping to ensure the trustworthiness of neural network classifier outputs. Given a neural network $N$ produces output class $\mathit{ArgMax}(N(x))$ for input $x$, we say that this answer is robust if the neural network would have outputted the same class for any nearby input $x^{\prime}$ within distance $\varepsilon$ of $x$: $\forall x^{\prime}.\ \|x-x^{\prime}\|\leq\varepsilon\Longrightarrow\mathit{ArgMax}(N(x^{\prime}))=\mathit{ArgMax}(N(x))$.
@@ 2.1. Robustness and Lipschitz-Based Certification
We write $B(x,\varepsilon):=\{x^{\prime}\mid\|x-x^{\prime}\|_{2}\leq\varepsilon\}$ for the closed $\varepsilon$-ball centred at $x$; robustness of $x$ is then the condition that $N$ classifies every $x^{\prime}\in B(x,\varepsilon)$ identically to $x$.
Fix an input $x$ and its associated output $y=N(x)$. Let $i^{*}=\mathit{ArgMax}(y)$ denote the output class of $x$, and let $j$ be any other competing class. We write $m_{j,i^{*}}(x)$ to denote the margin between the output and competiting class at input $x$, which is simply $y_{i^{*}}-y_{j}$. This definition generalises to arbitrary inputs $x^{\prime}$: $m_{j,i^{*}}(x^{\prime})=N(x^{\prime})_{i^{*}}-N(x^{\prime})_{j}$. Robustness of $x$ is enforced whenever this margin remains positive for all $x^{\prime}$ within $\varepsilon$ of $x$, i.e., whenever for each competitor class $j\not=i^{*}$ we have
@@
Methods for computing margin Lipschitz constants are outside the scope of this paper, but include iterative approximation algorithms, whether from below as in power iteration (Gouk et al., 2021), or from above as in Gram iteration (Delattre et al., 2023). With margin Lipschitz constants computed, certifying robustness for an input $x$ requires simply checking for each $j\not=i^{*}$ that
@@ 3.1. Feed-Forward Neural Networks
We begin by briefly recapping the standard model of these networks, over real arithmetic. Such networks are composed of $L$ layers. Let $\ell\in\{1,\ldots,L\}$ be a layer index. Then each layer $\ell$ comprises a weight matrix $W_{\ell}\in\mathbb{R}^{m_{\ell}\times n_{\ell}}$, bias vector $b_{\ell}\in\mathbb{R}^{m_{\ell}}$, and activation function $\varphi_{\ell}$. Let $x\in\mathbb{R}^{n_{1}}$ denote the neural network input and $z_{\ell-1}(x)$ denote the input to layer $\ell$ that is produced when $x$ is fed into the network. Then $z_{0}(x)=x$. Let $a_{\ell}$ denote the layer-$\ell$ preactivation, defined as
@@ 3.2. Floating-Point Error
The vector $\eta_{\mathrm{mv}}\in\mathbb{R}^{m}$ captures the accumulation of the per-product absolute terms across each length-$n$ dot product. A safe, input-independent bound is
@@
interpreted entrywise (here $\mathbbm{1}$ is the all-ones vector in $\mathbb{R}^{m}$). Passing to the $\ell_{2}$ norm, since each $|{\eta_{\mathrm{mv}}}_{i}|\leq a_{\mathrm{dot}}(n)$:
@@ 7. Pre-Deployment Hybrid Certification
In this section, we explore how these bounds can be made tighter in a pre-deployment setting, in which robustness certification is being used to evaluate a model before it is deployed. This is the setting in which robustness verifiers (Kaulen et al., 2025) are often applicable. Many such verifiers are orders of magnitude slower than robustness certifiers and so are more difficult to apply to evaluate model outputs at deployment-time, at which answers are needed in miliseconds. The pre-deployment setting is therefore an important use-case for robustness checkers.
Our key idea is that quantities like $D_{\ell}$ can be very conservative at lower-precision formats; however, are very tight at high-precision formats like float64. Also, while one’s goal might be to evalauate the robustness of a model executing in a standard format like float32, at pre-deployment time it is possible to take measurements from a high-precision execution of the model (e.g., float64) and to combine those with tight high-precision format deviation bounds to much more tightly bound deviation than the worst-case quantity $D_{\ell}$.
@@ 8.3.2. RQ3: Pre-Deployment Certification: Coverage and Competitiveness
The problem that our pre-deployment certification method (Section 7) solves is how to certify the robustness of a model over a distrubution, such as its standard test distribution. In this setting, the quantity of interest is the model’s Verified Robust Accuracy (VRA): the proportion of points that it correctly classifies and that are certified or verified to be robust. There are two questions of interest here: first, how much does FP-aware certification degrade VRA relative to a real arithmetic certifier? Second, how does our method compare to ERAN (Müller et al., 2018), which is—to our knowledge—the only publicly available robustness verifier that claims soundness against the floating-point execution semantics of neural networks and that has been applied to compute VRA statistics for models on the scale of those used in our evaluation.
@@ 9. Related Work
We motivated our investigation by constructing adversarial inputs that are misleadingly certified robust (often at very small $\varepsilon$) by Lipschitz-based certifiers, despite the existence of counterexample points within $\varepsilon$ that the model classifies differently due to floating-point rounding. Jia and Rinard (Jia and Rinard, 2021) first demonstrated a similar phenonenon for complete robustness verifiers, at much larger, semantically meaningful values of $\varepsilon$. Subsequent work also showed how to construct backdoored networks together with specific trigger inputs for which a complete verifier claims robustness under its arithmetic model, yet counterexamples exist under floating-point execution (Zombori et al., 2021). More recently, Szász et al. (Szász et al., 2025) go further and show how low-level details like the order of operations, mixed-precision implementations, and other specifics of the execution runtime, can lead to robustness verifiers giving misleading results for specially crafted networks. Our theory does not explicitly handle such details, if they depart from standard floating-point semantics. We return to this point in Section 10.
@@
Almost all of these tools focus on verification problems involving individual inputs, rather than producing distribution-level guarantees as is commmon for certified robustness mechanisms, such as Verified Robust Accuracy (VRA) which was the quantity of interest for our pre-deployment certification scenario studied in Section 8.3.2.
@@ 10. Conclusion
In this paper we identity how Lipschitz-based robustness certification can be semantically unsound under float-point neural network execution. To do so we showed concrete counterexamples that cause even a formally verified certifier to give misleading answers. Focusing on feed-forward networks with ReLU activations, we showwed how to certify the absence of overflow and constructed a compositional bound on the deviation between floating-point and real arithmetic execution, which can be operationalised through interleaved application layer-by-layer through a neural network. We showed how this bound can be applied by modifying the classical real arithmetic certification condition to account for floating-point rounding. This yields a quantification on the degradation of the robustness certificate under floating-point execution. We showed how this degradation can be halved when certification is performed pre-deployment, by taking advantage of deviation measurements from a high-precision (e.g., float64) execution. We mechanised our theory in the Rocq theorem prover, and implemented it within a Python based certifier. We empirically evaluated our certifier, demonstrating that our methods are practical for MNIST and CIFAR-10 scale networks.

### 17 Mar 2026 v2

@@ 1. Introduction
Robustness is an important property for helping to ensure the trustworthiness of neural network classifier outputs. Given a neural network $N$ produces output class $\mathit{ArgMax}(N(x))$ for input $x$, we say that this answer is robust if the neural network would have output the same class for any nearby input $x^{\prime}$ within distance $\varepsilon$ of $x$: $\forall x^{\prime}.\ \|x-x^{\prime}\|\leq\varepsilon\Longrightarrow\mathit{ArgMax}(N(x^{\prime}))=\mathit{ArgMax}(N(x))$.
@@ 2.1. Robustness and Lipschitz-Based Certification
We write $B(x,\varepsilon):=\{x^{\prime}\mid\|x-x^{\prime}\|_{2}\leq\varepsilon\}$ for the closed $\varepsilon$-ball centred at xx [^1] We write :=:= to mean “is defined as”.; robustness of $x$ is then the condition that $N$ classifies every $x^{\prime}\in B(x,\varepsilon)$ identically to $x$.
Fix an input $x$ and its associated output $y=N(x)$. Let $i^{*}=\mathit{ArgMax}(y)$ denote the output class of $x$, and let $j$ be any other competing class. We write $m_{j,i^{*}}(x)$ to denote the margin between the output and competing class at input $x$, which is simply $y_{i^{*}}-y_{j}$. This definition generalises to arbitrary inputs $x^{\prime}$: $m_{j,i^{*}}(x^{\prime})=N(x^{\prime})_{i^{*}}-N(x^{\prime})_{j}$. Robustness of $x$ is enforced whenever this margin remains positive for all $x^{\prime}$ within $\varepsilon$ of $x$, i.e., whenever for each competitor class $j\not=i^{*}$ we have
@@
Methods for computing margin Lipschitz constants are outside the scope of this paper. Briefly, they are derived from the spectral norms of the layer weight matrices. These spectral norms are computed by iterative approximation algorithms, whether from below as in power iteration (Gouk et al., 2021), or from above as in Gram iteration (Delattre et al., 2023). With margin Lipschitz constants computed, certifying robustness for an input $x$ requires simply checking for each $j\not=i^{*}$ that
@@ 3.1. Feed-Forward Neural Networks
We begin by briefly recapping the standard model of these networks, over real arithmetic. Such networks are composed of $L$ layers. Let $\ell\in\{1,\ldots,L\}$ be a layer index. Then each layer $\ell$ comprises a weight matrix $W_{\ell}\in\mathbb{R}^{m_{\ell}\times n_{\ell}}$, bias vector $b_{\ell}\in\mathbb{R}^{m_{\ell}}$, and activation function $\varphi_{\ell}$. Layer $\ell$ transforms an $n_{\ell}$-element input vector (the output of layer $\ell-1$) into an $m_{\ell}$-element output vector (the input to layer $\ell+1$). Let $x\in\mathbb{R}^{n_{1}}$ denote the neural network input and $z_{\ell-1}(x)$ denote the input to layer $\ell$ that is produced when $x$ is fed into the network. Then $z_{0}(x)=x$. Let $a_{\ell}$ denote the layer-$\ell$ preactivation, defined as
@@ 3.2. Floating-Point Error
Here $\Delta{W}$ is a matrix whose entries depend on $W$ and $z$. For a matrix $M$ or vector $v$, we write $|M|$ or $|v|$ to denote the matrix or vector formed by taking the absolute value entrywise. We overload $\leq$ on matrices and vectors, where it is interpreted entrywise. The vector $\eta_{\mathrm{mv}}\in\mathbb{R}^{m}$ captures the accumulation of the per-product absolute error terms across each length-$n$ dot product. A safe, input-independent bound is
@@
where $\mathbbm{1}$ is the all-ones vector in $\mathbb{R}^{m}$. Passing to the $\ell_{2}$ norm, since each $|{\eta_{\mathrm{mv}}}_{i}|\leq a_{\mathrm{dot}}(n)$:
@@ 7. Pre-Deployment Hybrid Certification
In this section, we explore how these bounds can be made tighter in a pre-deployment setting, in which robustness certification is being used to evaluate a model before it is deployed. This is the setting in which robustness verifiers (Kaulen et al., 2025) are often applicable. Many such verifiers are orders of magnitude slower than robustness certifiers and so are more difficult to apply to evaluate model outputs at deployment-time, at which answers are needed in milliseconds. The pre-deployment setting is therefore an important use-case for robustness checkers.
Our key idea is that quantities like $D_{\ell}$ can be very conservative at lower-precision formats; however, are very tight at high-precision formats like float64. Also, while one’s goal might be to evaluate the robustness of a model executing in a standard format like float32, at pre-deployment time it is possible to take measurements from a high-precision execution of the model (e.g., float64) and to combine those with tight high-precision format deviation bounds to much more tightly bound deviation than the worst-case quantity $D_{\ell}$.
@@ 8.3.2. RQ3: Pre-Deployment Certification: Coverage and Competitiveness
The problem that our pre-deployment certification method (Section 7) solves is how to certify the robustness of a model over a distribution, such as its standard test distribution. In this setting, the quantity of interest is the model’s Verified Robust Accuracy (VRA): the proportion of points that it correctly classifies and that are certified or verified to be robust. There are two questions of interest here: first, how much does FP-aware certification degrade VRA relative to a real arithmetic certifier? Second, how does our method compare to ERAN (Müller et al., 2018), which is—to our knowledge—the only publicly available robustness verifier that claims soundness against the floating-point execution semantics of neural networks and that has been applied to compute VRA statistics for models on the scale of those used in our evaluation.
@@ 9. Related Work
We motivated our investigation by constructing adversarial inputs that are misleadingly certified robust (often at very small $\varepsilon$) by Lipschitz-based certifiers, despite the existence of counterexample points within $\varepsilon$ that the model classifies differently due to floating-point rounding. Jia and Rinard (Jia and Rinard, 2021) first demonstrated a similar phenomenon for complete robustness verifiers, at much larger, semantically meaningful values of $\varepsilon$. Subsequent work also showed how to construct backdoored networks together with specific trigger inputs for which a complete verifier claims robustness under its arithmetic model, yet counterexamples exist under floating-point execution (Zombori et al., 2021). More recently, Szász et al. (Szász et al., 2025) go further and show how low-level details like the order of operations, mixed-precision implementations, and other specifics of the execution runtime, can lead to robustness verifiers giving misleading results for specially crafted networks. Our theory does not explicitly handle such details, if they depart from standard floating-point semantics. We return to this point in Section 10.
@@
Almost all of these tools focus on verification problems involving individual inputs, rather than producing distribution-level guarantees as is common for certified robustness mechanisms, such as Verified Robust Accuracy (VRA) which was the quantity of interest for our pre-deployment certification scenario studied in Section 8.3.2.
@@ 10. Conclusion
In this paper we identify how Lipschitz-based robustness certification can be semantically unsound under floating-point neural network execution. To do so we showed concrete counterexamples that cause even a formally verified certifier to give misleading answers. Focusing on feed-forward networks with ReLU activations, we showed how to certify the absence of overflow and constructed a compositional bound on the deviation between floating-point and real arithmetic execution, which can be operationalised through interleaved application layer-by-layer through a neural network. We showed how this bound can be applied by modifying the classical real arithmetic certification condition to account for floating-point rounding. This yields a quantification on the degradation of the robustness certificate under floating-point execution. We showed how this degradation can be halved when certification is performed pre-deployment, by taking advantage of deviation measurements from a high-precision (e.g., float64) execution. We mechanised our theory in the Rocq theorem prover, and implemented it within a Python based certifier. We empirically evaluated our certifier, demonstrating that our methods are practical for MNIST and CIFAR-10 scale networks.
Our results also shed light on the practical significance of the semantic gap between real arithmetic certification and floating-point execution. Although Lipschitz-based certification is conservative—using input-independent bounds on sensitivity that over-estimate the model’s local behaviour—our counterexamples (Section 2.3) show that this conservatism alone does not mask floating-point discrepancies—even if the observed discrepancies are small at standard precisions (e.g., float32 and above). Moreover, our floating-point-aware analysis currently has limited applicability in the lowest-precision regimes (e.g., float16) where floating-point discrepancies become largest (Section 8). These observations clarify when floating-point effects are most likely to matter for Lipschitz-based robustness certification, while our theory and implementation provide the first method for soundly accounting for these effects in Lipschitz-based robustness certification.